# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ Abstract
This paper presents the Holistic and Multi-Granular Surgical Scene Understanding of Prostatectomies (GraSP) dataset, a curated benchmark that models surgical scene understanding as a hierarchy of complementary tasks with varying levels of granularity. Our approach enables a multi-level comprehension of surgical activities, encompassing long-term tasks such as surgical phases and steps recognition and short-term tasks including surgical instrument segmentation and atomic visual actions detection. To exploit our proposed benchmark, we introduce the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model, a general architecture that combines a global video feature extractor with localized region proposals from an instrument segmentation model to tackle the multi-granularity of our benchmark. Through extensive experimentation, we demonstrate the impact of including segmentation annotations in short-term recognition tasks, highlight the varying granularity requirements of each task, and establish TAPIS’s superiority over previously proposed baselines and conventional CNN-based models. Additionally, we validate the robustness of our method across multiple public benchmarks, confirming the reliability and applicability of our dataset. This work represents a significant step forward in Endoscopic Vision, offering a novel and comprehensive framework for future research towards a holistic understanding of surgical procedures.
@@ 1 Introduction
Robot-assisted surgery (RAS) has emerged as a pivotal advancement in modern surgical practices (Goh and Ali (2022)), offering a minimally invasive option compared to common open surgeries and providing enhanced precision over conventional laparoscopic techniques (Chuchulo and Ali (2023)). This innovative approach provides enhanced instrumentation, visualization, and dexterity capabilities, greatly benefiting surgeons in the operating room and significantly reducing patient complications (Lanfranco et al. (2004); van Amsterdam et al. (2022)). However, robotic surgery currently faces several limitations, such as the lack of automated assistance from the robot (Fiorini (2021)), minimal image-based force feedback (Rassi and Rassi (2020)), and a steep learning curve for specific surgeries (Kim et al. (2019); Vining et al. (2021)). Nevertheless, the extensive and complex data collected by surgical robots provide a unique opportunity to develop data-based solutions that enrich the robot with surgical intelligence (Maier-Hein et al. (2017)). These advancements would augment surgical processes, offering ways to overcome multiple challenges in robotic surgical practices.
A step towards exploiting the full potential of surgical robotic systems is the comprehension of the visual information relayed by the robot’s endoscope (Mascagni et al. (2022)). This understanding extends beyond merely recognizing and categorizing present elements, such as anatomical structures and surgical instruments, in the visual data. Instead, it involves fully integrating the temporal and spatial dimensions for a coherent and holistic interpretation of surgical scenes that can be useful for the robotic system. Achieving this higher level of understanding can lead to enhanced augmented reality systems (Qian et al. (2020); Tanzi et al. (2021)), context-aware surgical assistance (Katić et al. (2013); Kolbinger et al. (2023)), improved surgical education methods (Maier-Hein et al. (2017)), and even the development of semi-automated surgical interventions.
In light of this, Endoscopic Vision is a critical area of Surgical Data Science that concentrates on interpreting and analyzing the rich visual information captured during endoscopic procedures (Fu et al. (2021); Maier-Hein et al. (2022)). This field tackles surgical scene understanding with multiple methodological approaches representing essential components of image-guided surgical cognition like comprehension of environmental geometry (Allan et al. (2020); Wang et al. (2022)), surgical procedure knowledge (Maier-Hein et al. (2021); Wagner et al. (2023)), and understanding of agent-environment relations (Islam et al. (2020); Nwoye et al. (2020)). The specific visual cues that these approaches prioritize or the particular aspects of the surgical environment they study lead to distinct requirements in terms of visual recognition, temporal awareness, and spatial granularity. Among these approaches, surgical workflow analysis studies the complex interactions between surgical instruments and anatomical landmarks (Padoy et al. (2012); Valderrama et al. (2022)). This area aims to comprehend the temporal progression of surgical procedures by understanding the activities performed throughout the sequence of surgical scenes in a surgical video. For instance, the most studied task in surgical workflow analysis is the recognition of surgical phases in a given endoscopic video clip or image (Demir et al. (2023)). Even so, surgical workflow analysis has an implicit need for a multi-recognition approach that identifies the different agents and activities in the surgical setting (Twinanda et al. (2016); Maier-Hein et al. (2021)). Hence, other surgical workflow approaches extend to the recognition of surgical steps, surgical gestures, the presence of surgical instruments, or the presence of actions (Demir et al. (2023)).
@@
Despite the advancements in multi-task approaches of surgical workflow analysis benchmarks, there remains a significant gap in the literature concerning a holistic understanding of the multi-level nature of surgical interventions. A fundamental limitation of most current surgical workflow analysis benchmarks is their constrained study of the spatial dimensions of endoscopic images. Specifically, current benchmarks often disregard the visual localization of the constitutive acting agents, which are the surgical instruments in a surgical procedure (Nwoye et al. (2022)). Notably, just a few benchmarks of laparoscopic surgeries integrate surgical instrument detection or segmentation into surgical workflow analysis, a gap primarily due to the time-consuming nature of gathering these annotation types (Maier-Hein et al. (2021); Grammatikopoulou et al. (2021)). This oversight is evident, as instrument segmentation, detection, or tracking, despite their critical relevance, have mostly been treated as independent areas from surgical workflow analysis. For example, the Endoscopic Vision challenges on instrument segmentation (Bodenstedt et al. (2018); Allan et al. (2019, 2020)) have become the standard benchmarks on pixel-wise instrument recognition, but besides external annotations of instrument-tissue interactions (Islam et al. (2020)), these frameworks do not provide annotations for other surgical workflow analysis tasks.
Nonetheless, integrating visual locations of surgical instruments has primordial implications for robot-assisted surgery, as current robotic systems generate instrument utilization signals but lack the capability to relate these presence data to their visual locations within the surgical scene (Du et al. (2016); Su et al. (2018); Guan et al. (2023)). On top of that, instrument segmentation would allow surgical robotic systems to benefit from the most granular approach to instrument recognition and thus acquire a notion of instrument shape (Baby et al. (2023)), pose (Du et al. (2018)), and pixel-wise boundaries (García-Peraza-Herrera et al. (2017)), which, in turn, can lead to multiple augmented reality applications (Qian et al. (2020); De Backer et al. (2023)). Therefore, bridging this gap between instrument segmentation and surgical workflow analysis is essential for advancing the field of Robotic Endoscopic Vision.
A second area for improvement in current surgical workflow analysis and Endoscopic Vision benchmarks, closely related to the first, concerns the exploitation of the hierarchical complementarity among different visual perception tasks in Endoscopic Vision. Initially, some benchmarks treat recognition tasks such as phases recognition (Stauder et al. (2017)) and action recognition (Bawa et al. (2021); Goodman et al. (2021)) as isolated entities rather than interdependent components of a cohesive surgical process. For example, as shown in Table 1, the LapChole dataset (Stauder et al. (2017)) focuses on the sole prediction of surgical phases in traditional Laparoscopic Surgeries (LS), the CATARACTS dataset (Al Hajj et al. (2019)) initially focused only on tool presence recognition in Cataract Surgery (CS), and the JIGSAWS (Ahmidi et al. (2017)), LRSAO (Sharghi et al. (2020)), ESAD (Bawa et al. (2021)) and RARP45 (van Amsterdam et al. (2022)) benchmarks only study gestures and actions in either RAS data or Robot-Assisted Surgical Training (ST) data.
Similarly, multi-task benchmarks tend to focus on a few levels of visual understanding, like frame-level binary presence prediction or video-level temporal segmentation of phases, rather than tackling a comprehensive range of levels and granularities in the visual, temporal, and spatial dimensions. The Cholec80 benchmark (Twinanda et al. (2016)) and the re-annotated CATARACTS benchmark (Al Hajj et al. (2019); Zisimopoulos et al. (2018)) study one video-level task (phase recognition) with one frame-level presence recognition task (instrument presence) in LS data. The CholecT40 (Nwoye et al. (2020)), CholecT50 (Nwoye et al. (2022)), and HeiChole (Wagner et al. (2023)) benchmarks include more frame-level presence recognition tasks like the presence of actions and anatomical targets in LS as well. Alternatively, the MISAW framework (Huaulmé et al. (2021)) is the only dataset to consider multiple granularities in video-level tasks by annotating phases, steps, and activities in ST data. Conversely, the HeiCo (Maier-Hein et al. (2021)) and CaDIS (Grammatikopoulou et al. (2021)) datasets are pioneering benchmarks in combining video-level phase recognition with pixel-wise instrument recognition in LS and CS data, but they disregard action information. Still, existing frameworks predominantly focus on single video-level tasks with some frame-wise recognition tasks without adequately addressing the multiple hierarchical granularities required for holistic surgical workflow analysis.
@@
In our initial work (Valderrama et al. (2022)), we laid the foundation for studying surgical scenes across multiple levels of granularity in both the temporal and spatial dimensions. The empirical framework presented herein models surgical procedures as a hierarchy of semantic tasks encompassing long-term video-level tasks, such as phase and step recognition, and short-term spatio-temporal tasks, including instrument and atomic visual action localization. We base our modeling on the Atomic Visual Actions (AVA) formulation in Gu et al. (2018) and the activity understanding theory of Barker and Wright (1954). The latter presents a hierarchical comprehension of activity, where the finest expressions of activity are the atomic interactions and movements performed by the main acting agents of a scene, and the highest comprehension level of activity corresponds to the development of complex and general objectives. Our work adapts these concepts into surgical workflow analysis. Long-term recognition of phases and steps corresponds to a higher-level understanding of sequences of necessary goals in surgical reasoning. In contrast, short-term tasks of identifying instruments and their atomic actions are akin to the lower-level notion of atomic movements done by an acting agent.
Thus, the rationale behind our framework hinges on a stratification of surgical analysis tasks where the apex is the recognition of surgical phases, representing the highest and most general level of understanding. These phases can be further decomposed into detailed steps, a more fine-grained level of reasoning. Each step involves the utilization of particular surgical instruments, which perform specific atomic actions. We introduce surgical atomic actions as our framework’s lowest and most fine-grained level of understanding, which provides a practical way of studying surgical activities and allows the further reconstruction of the activity hierarchy (Gu et al. (2018)). By breaking down the procedure into these distinct yet interconnected levels, we emphasize the importance of understanding the individual components and their collective dynamics. Such analysis is crucial for a holistic understanding of the surgical scene, capturing the essence of both the macro and micro-elements of surgeries.
@@
In this work, we extend our original framework (Valderrama et al. (2022)) by integrating pixel-wise recognition of instruments to achieve the highest level of spatial recognition granularity. For this purpose, we transformed our original instrument detection annotations into detailed instrument segmentation annotations. Hence, we introduce a new instrument segmentation task in our dataset and a fully instance-based evaluation benchmark that aligns more closely with the practical needs of surgical scene analysis. Additionally, we augment our dataset with five new fully annotated videos, and we have established specific training, validation, and testing splits with per-task metrics for benchmarking. The newly annotated videos form the test set, while the previous videos are the training set divided into two explicitly defined folds for a two-fold cross-validation process. We curate and revise all our annotations to eliminate noise and inconsistencies, ensuring the highest level of data integrity. Our extension in data accounts for more than a 55% increase in data size and annotation over the initial PSI-AVA dataset. This advanced version of our framework is called the Holistic and Multi-Granular Surgical Scene Understanding of Prostatectomies (GraSP) dataset. Table 1 shows the contributions of our benchmark compared to previous literature. The discussed improvements make our benchmark the first publicly available in the robotic surgery domain with a multi-task approach and an instrument instance segmentation task using real in-patient surgical data on human subjects. To our knowledge, GraSP is the first Endoscopic Vision framework to consider multiple long and short-term tasks in their most fine-grained form.
@@
Besides the extensions to our benchmark, we propose the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model. A generalized, fully transformer-based architecture that combines a global video feature extractor with a localized region proposal network to tackle each task in the GraSP benchmark. TAPIS builds upon our previous outstanding methodologies for holistic surgical scene understanding (Valderrama et al. (2022)) and instrument segmentation (Ayobi et al. (2023)). To fully include our novel instrument segmentation benchmark, we transition from the instrument detector in Valderrama et al. (2022) into an instrument segmentation baseline that provides detailed region proposals and shape-wise features of surgical instruments. Our experimentation demonstrates that these regions and their features consistently enhance recognition accuracy and enrich the atomic action recognition capabilities. We also propose an improved region classification head that exploits transformers’ cross-attention to fully incorporate the totality of the rich temporal embeddings computed by the video feature extractor. Finally, we validate our method’s robustness and applicability by comparing it with alternative baselines and our previous approaches and testing it in various publicly available frameworks. Our results demonstrate that TAPIS achieves outstanding performances in multiple datasets and establishes the state-of-the-art in our GraSP benchmark. Our efforts in dataset extension, methodological innovation, and benchmark validation significantly contribute to the research in surgical data science and pave the way towards a holistic understanding of surgical procedures.
@@
    We propose the GraSP dataset for holistic surgical scene understanding, a curated and augmented version of the PSI-AVA benchmark, with defined benchmark splits and more than a 55% increase in data and annotations.
@@
    We introduce a new instrument segmentation task into our benchmark as the most granular approach to instrument localization.
@@
    We present a generalized, transformer-based architecture that strongly outperforms all previous baselines across all tasks of the GraSP framework.
@@ 2 Related Work
Table 1: Our GraSP benchmark compared with current surgical workflow analysis and endoscopic vision datasets. Each row represents a current Endoscopic Vision dataset, and each column is a possible dataset attribute corresponding to either the inclusion of multiple levels of semantic comprehension, annotations of a particular surgical scene understanding the task, video duration, public availability, or data source type. We show whether an attribute is present (✓), absent (–), has a different formulation (*), or its information is unavailable (?). We also indicate if the source of the provided data is from traditional Laparoscopic Surgery (LS), ex vivo Surgical Training Procedures (ST), Cataract Surgery (CS), Robot-Assisted Surgery on animal subjects (RASA), Robot-Assisted Surgery on human subjects (RASH), or multiple Open Surgery types (MOS). Our GraSP dataset, shown in bold, is the first publicly available benchmark to provide all the presented attributes. Note: We combine the Instrument Detection and Instrument Segmentation attributes to identify datasets with instance-level annotations. Datasets with both attributes have Instrument Instance Segmentation annotations, datasets with only Instrument Detection have bounding box annotations, and datasets with only Instrument Segmentation annotations have Semantic Segmentation annotations.
@@
Cholec80 (Twinanda et al. (2016)) 51.25 LS
TUM LapChole (Stauder et al. (2017)) 23.24 LS
JIGSAWS (Ahmidi et al. (2017)) ✓* * 2.62 ST
Nephrec9 (Ahmidi et al. (2017)) 10.52 RASH
EndoVis 2015 (Bodenstedt et al. (2018)) 0.10 LS, ST
EndoVis 2017 (Allan et al. (2019)) 0.83 RASA
CATARACTS (Al Hajj et al. (2019); Zisimopoulos et al. (2018)) 9.2 CS
CaDIS (Zisimopoulos et al. (2018); Grammatikopoulou et al. (2021)) 9.1 CS
EndoVis 2018 (Allan et al. (2020); González et al. (2020); Islam et al. (2020)) 1.58 RASA
LSRAO (Sharghi et al. (2020)) ? RASH
CholecT40, T45 & T50 (Nwoye et al. (2020, 2022)) 28.03 LS
HeiCo (Maier-Hein et al. (2021); Roß et al. (2021)) 96.12 LS
MISAW (Huaulmé et al. (2021)) –* 1.52 ST
ESAD (Bawa et al. (2021)) 9.33 RASH
AVOS (Goodman et al. (2021)) 47 MOS
RARP45 (van Amsterdam et al. (2022)) ✓* 3.2 RASH
RARR (Kolbinger et al. (2023)) 393.3 RASH
HeiChole (Wagner et al. (2023)) 22 LS
PSI-AVA (Valderrama et al. (2022)) 20.45 RASH

### v3

@@ Abstract
This paper presents the Holistic and Multi-Granular Surgical Scene Understanding of Prostatectomies (GraSP) dataset, a curated benchmark that models surgical scene understanding as a hierarchy of complementary tasks with varying levels of granularity. Our approach encompasses long-term tasks, such as surgical phase and step recognition, and short-term tasks, including surgical instrument segmentation and atomic visual actions detection. To exploit our proposed benchmark, we introduce the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model, a general architecture that combines a global video feature extractor with localized region proposals from an instrument segmentation model to tackle the multi-granularity of our benchmark. Through extensive experimentation in ours and alternative benchmarks, we demonstrate TAPIS’s versatility and state-of-the-art performance across different tasks. This work represents a foundational step forward in Endoscopic Vision, offering a novel framework for future research towards holistic surgical scene understanding.
@@ 1 Introduction
Robot-assisted surgery (RAS) has emerged as a major advancement in modern surgical practices ([31]), offering a minimally invasive alternative to open surgeries and providing enhanced precision over conventional laparoscopic techniques ([14]). Its enhanced instrumentation, visualization, and dexterity greatly benefit surgeons and significantly reduce patient complications ([49, 4]). Despite these benefits, robotic surgery faces several limitations, including limited automated assistance ([26]), minimal image-based force feedback ([70]), and a steep learning curve for specific surgeries ([46, 86]). Nevertheless, the extensive data generated by surgical robots provide opportunities to develop data-based solutions that enrich the robot with surgical cognition and overcome current challenges ([59]).
A crucial step towards exploiting the potential of surgical robotic systems is understanding the visual data captured by the robot’s endoscope ([61]). Albeit, to fully interpret endoscopic data, it is necessary to go beyond simple object identification and adopt a holistic approach that thoroughly models the temporal and spatial dimensions of surgical videos. Such advanced understanding would enhance robotic systems by enabling augmented reality systems ([69, 81]), context-aware assistance ([43, 47]), improved surgical training ([59]), and development of semi-automated surgical interventions.
In light of this, Endoscopic Vision (EV) is the area within Surgical Data Science that studies the essential components of image-guided surgical cognition like surgical environment geometry ([88]), surgical procedural knowledge ([87]), and tools-tissues interactions ([39, 63]). However, each methodological approach in EV prioritizes different visual cues that lead to different requirements in visual recognition, temporal awareness, and spatial granularity. For example, surgical workflow analysis (SWA) studies tasks such as surgical phase recognition to understand the temporal progression of surgical procedures ([66]). Nevertheless, this objective implicitly requires a multirecognition methodology that identifies different agents and activities over time in surgical settings ([83, 65]). Consequently, SWA has expanded to multitask approaches that include recognizing surgical steps, gestures, actions, or instruments  ([17]).
@@
Despite the advancements in multi-task benchmarks for SWA, there remains a significant gap in the literature concerning a holistic understanding of the multi-level nature of surgical interventions. A fundamental limitation of most current SWA benchmarks is their constrained study of the spatial dimensions of endoscopic images. Specifically, current benchmarks often disregard the visual localization of surgical instruments, which are the constitutive acting agents in any endoscopic surgery ([65]). Instead, surgical instrument segmentation or detection tasks have mostly been treated as independent areas from SWA ([40]), and very few endoscopic benchmarks integrate these tasks with SWA due to the time-consuming nature of gathering such annotations. Nonetheless, integrating the visual locations of surgical instruments is primordial for robot-assisted surgery, as current robotic systems can track instrument usage and kinematics but cannot link these signals to the instruments’ visual locations ([21, 80, 35]). On top of that, instrument segmentation, as the finest-grained level of visual recognition, would allow surgical robots to understand instrument shapes ([6]), poses ([22]), and visual boundaries ([30]).
A second area for improvement in current SWA benchmarks concerns the exploitation of the hierarchical complementarity among different visual perception tasks. Initially, some benchmarks treat single recognition tasks such as phases recognition ([79]) or action recognition ([8, 33]) as isolated tasks rather than interdependent components of a cohesive surgical process. Similarly, multi-task benchmarks focus on a few levels of visual understanding, like one video-level task with several frame-level presence tasks ([83, 65]), or multiple video-level tasks with no spatial localization ([36]), rather than tackling multiple levels in the visual, temporal, and spatial dimensions.
@@
In our initial work ([84]), we laid the foundation for studying surgical scenes across multiple temporal and spatial granularities by modeling surgical procedures as a hierarchy of long-term video-level tasks, such as phase and step recognition, and short-term spatio-temporal tasks including instrument and atomic visual action localization. We base our modeling on the Atomic Visual Actions formulation in [34] and the activity theory of [7], which define activity as a hierarchy starting from its finest expression as atomic interactions performed by acting agents, to the broader development of complex objectives. Our work adapts these concepts into surgical workflow analysis with a stratification of semantic tasks where surgical phase recognition represents the most general level of understanding. These phases can be further decomposed into finer-grained steps, where each step involves performing specific atomic actions that require using particular surgical instruments. We introduced surgical atomic actions as a spatially localized approach to surgical action recognition and the finest temporal modeling of surgical activities. We portray this hierarchical structure in Figure 2. By breaking down the procedure into these distinct yet interconnected levels, we emphasize the importance of understanding the individual components and their collective dynamics, capturing the essence of both the macro and micro-elements of surgeries.
@@
In this work, we extend our original framework by integrating an instrument instance segmentation task to achieve the highest level of spatial granularity. Additionally, we augment our dataset with five new fully annotated videos, curate all our annotations, and define specific data splits and metrics for benchmarking. Our extension in data accounts for more than a 55% increase in data size and annotation over the initial PSI-AVA dataset ([84]). We call this advanced framework the Holistic and Multi-Granular Surgical Scene Understanding of Prostatectomies (GraSP) dataset. These improvements make our benchmark the first publicly available in the robotic surgery domain to include phase and step annotations and an instrument instance segmentation task using real human in-patient surgical data. To our knowledge, GraSP is the first EV framework to consider multiple long and short-term tasks in their most fine-grained form.
@@
Besides the extensions to our benchmark, we propose the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model. We introduce TAPIS as a versatile, fully transformer-based architecture designed to tackle each task in the GraSP benchmark. Its flexibility allows it to function as both a multi-task framework, leveraging complementary tasks with varying granularities, or as a modular system capable of handling disjoint tasks independently. This adaptability enables TAPIS to seamlessly adapt to diverse endoscopic vision datasets and accommodate a wide range of video recognition tasks. TAPIS builds upon our previous methodologies for holistic surgical scene understanding ([84]) and segmentation ([5]) by incorporating an instrument segmentation baseline that provides region proposals and shape-wise region features that enhance instrument and action recognition. We also introduce an improved region classification head that exploits transformers’ cross-attention to fully integrate temporal and spatial embeddings. Through comparisons with alternative baselines and SOTA models on GraSP and publicly available benchmarks, we demonstrate that TAPIS consistently achieves state-of-the-art performance across diverse tasks. Our efforts in dataset extension, methodological innovation, and benchmark validation, contribute to surgical data science by paving the way towards a holistic understanding of surgical procedures.
@@
    We propose the GraSP dataset for holistic surgical scene understanding, a curated and augmented version of the PSI-AVA benchmark with defined benchmark splits.
@@
    We introduce a new instrument instance segmentation task into our benchmark.
@@
    We present a generalized, transformer-based architecture that strongly outperforms alternative baselines across all tasks of the GraSP framework.
@@
Fig. 2: Hierarchical Structure of GraSP Tasks. This figure illustrates the hierarchical organization of tasks in GraSP based on temporal and spatial granularity. Nodes represent category labels within each task level, while edges denote connections between categories across levels. The intensity of the edge color reflects the co-occurrence of the linked categories based on annotations from GraSP. To enhance visualization, we reorganized the categories and removed weakly connected edges, as well as highly persistent action categories.
@@ 2 Related Work
Table 1: Our GraSP benchmark compared with current surgical workflow analysis and endoscopic vision datasets. Each row represents a current Endoscopic Vision dataset, and each column is a possible dataset attribute corresponding to either the inclusion of multiple levels of semantic comprehension, annotations of a particular surgical scene understanding the task, video duration, public availability, or data source type. We show whether an attribute is present (✓), absent (–), has a different formulation (*), or its information is unavailable (?). We also indicate if the source of the provided data is from traditional Laparoscopic Surgery (LS), ex vivo Surgical Training Procedures (ST), Robot-Assisted Surgery on animal subjects (RASA), or Robot-Assisted Surgery on human subjects (RASH). Our GraSP dataset, shown in bold, is the first publicly available benchmark to provide all the presented attributes. Note: We combine the Instrument Detection and Instrument Segmentation attributes to identify datasets with instance-level annotations. Datasets with both attributes have Instrument Instance Segmentation annotations, datasets with only Instrument Detection have bounding box annotations, and datasets with only Instrument Segmentation annotations have Semantic Segmentation annotations.
@@
Cholec80 ([83]) 51.25 LS
M2CAI16 ([79]) 23.24 LS
JIGSAWS ([1]) * 2.62 ST
Nephrec9 ([1]) 10.52 RASH
EndoVis 2015 ([9]) 0.10 LS, ST
EndoVis 2017 ([3]) 0.83 RASA
EndoVis 2018 ([2, 32, 39]) 1.58 RASA
LSRAO ([75]) ? RASH
CholecT40, T45 & T50 ([63, 65]) 28.03 LS
HeiCo ([60, 72]) 96.12 LS
MISAW ([36]) * 1.52 ST
ESAD ([8]) 9.33 RASH
SAR-RARP50 ([4, 68]) 3.2 RASH
AutoLaparo ([89]) 23.13 RASH
RARR ([47]) 393.3 RASH
HeiChole ([87]) 22 LS
PSI-AVA ([84]) 20.45 RASH